OpenStudy (anonymous):

find the approximate solution 5e^x =4

OpenStudy (nurali):

5e^x =4 e^x=4/5 e^x=0.8 Multiply Ln on both side lne^x=ln(0.8) xln(e)=-0.233 ln(e)=1 x=-0.233

OpenStudy (anonymous):

the answer is x=-0.2231 (rounded in four decimal places

OpenStudy (nurali):

Yes